# SuPreME: A Supervised Pre-training Framework for Multimodal ECG Representation Learning

### v1

@@ 3.3 Zero-shot Prompted Classification
Cardiac Query Design. In our framework, zero-shot classification is achieved by converting cardiac conditions in unseen downstream datasets into descriptive text queries. All SCP codes from the downstream datasets are compiled and paired with concise, clinically relevant textual descriptions, which serve as inputs to the textual modality of SuPreME. These cardiac queries enable the model to effectively align previously unseen cardiac conditions with ECG signal features, allowing it to generalize beyond the entities encountered during pre-training and facilitating inference on novel categories within downstream datasets.
To construct these cardiac queries, we employ simplified Clinical Knowledge-Enhanced Prompt Engineering (CKEPE) (Liu et al., 2024). While traditional CKEPE involves querying LLMs to extract detailed clinical knowledge from external medical databases (i.e., UMLS and SNOMED CT), including disease subtypes and signal pattern attributes, our approach follows a similar pipeline but directly interprets SCP codes from downstream datasets into concise cardiac descriptions. The robust representation capabilities of our CFN allow us to rely on simplified query content without compromising performance. By reducing redundancy in the queries, this method avoids overloaded textual information that might hinder the model’s ability to align features effectively across modalities.
Evaluation with Pre-trained SuPreME. During the evaluation process, ECGs and cardiac queries are encoded into $\mathbf{Emb}_{\text{ECGs}}$ and $\mathbf{Emb}_{\text{Queries}}$ respectively, using the ECG encoder and text encoder components of the pre-trained SuPreME framework. The embeddings are then input into the pre-trained CFN for alignment and fusion, which processes them using cross-modal attention mechanisms and outputs logits through an MLP head (i.e., a linear classification layer). Subsequently, the logits are passed through a sigmoid activation function $\sigma$ to produce probabilities $\mathbf{Pred}$ for each cardiac condition:
$\mathbf{Pred}=\sigma(\text{CFN}(\mathbf{Emb}_{\text{ECGs}},\mathbf{Emb}_{\text{Queries}}))$ (4)
We assess the model’s classification performance using the Area Under the Receiver Operating Characteristic curve (AUROC, also referred to as AUC; further details are provided in Appendix A.3). Given the multi-label nature of this classification task, the AUC is computed separately for each category, and the mean AUC across all categories is reported as the overall performance metric.
Overlap Analysis on Pre-train and Downstream Datasets. We analyze the cardiac query overlap between the pre-train dataset and six downstream datasets specified in Section 4.1, as well as among the downstream datasets themselves. Specifically, we embed all entities from the pre-train dataset and cardiac queries from the downstream datasets, compute their cosine similarity, and apply a threshold of 0.95 to filter overlapping queries. This analysis reveals that 57 cardiac queries overlap between the pre-train dataset and the downstream datasets. Details of the overlapping queries are provided in the Appendix A.4.
Figure 7: Overlap between pre-train and downstream datasets, with left panel showing pairwise downstream overlaps and right panel showing pre-train overlaps.
Furthermore, we analyze the overlaps both among the downstream datasets and between each downstream dataset and the pre-train dataset, as illustrated in Figure 7. The heatmap on the left shows that pairwise overlaps among downstream datasets are generally limited, reflecting the diversity of cardiac query prompts. The bar chart on the right shows that while the pre-train dataset shares some similar queries, a substantial portion of queries remains unique to the downstream datasets, allowing the pre-train process to establish robust general-purpose representations while leaving room for downstream-specific adaptation.
4 Experiments
4.1 Configuration and Settings
Clinical Entity Extraction. We utilize domain-specific prompts enriched with professional medical knowledge to extract entities from the MIMIC-IV-ECG dataset (see Appendix A.5). Entity extraction is performed using Llama3.1-70B-Instruct with structured constraints for clinical NER. Deduplication and mapping leverage MedCPT embeddings to group similar entities (similarity > 0.8) and map them to standardized terminologies from UMLS and SNOMED CT (average similarity > 0.75). Experiments are conducted on 8 NVIDIA A100-SMX4-80GB GPUs with vLLM [^6] https://github.com/vllm-project/vllm..
Supervised ECG Pre-training. A 1D ViT-tiny model serves as the ECG encoder, with a patch size of 125 (0.25 seconds per patch). MedCPT, with frozen weights (Zhang et al., 2023b), is used as the text encoder. We use the AdamW optimizer with a learning rate of $1\times 10^{-3}$ and weight decay of $1\times 10^{-8}$. A cosine annealing scheduler adjusts the learning rate with an initial cycle length $T_{0}$ of 5000 steps and $T_{\text{mult}}=1$, down to a minimum learning rate of $1\times 10^{-8}$. Pre-training runs for 100 epochs with early stopping tolerance as 10 epochs and uses a batch size of 256 per GPU on 4 NVIDIA A100-PCIE-40GB GPUs.
Downstream Classification Task. Zero-shot classification is evaluated on six unseen datasets (e.g., PTB-XL, CPSC-2018, and Chapman-Shaoxing-Ningbo; see Appendix A.5) using SuPreME and customized prompts created via CKEPE (Section 3.3). Ablation studies assess the impact of different ECG and text encoder backbones and the CFN module. Mainstream eSSLs are evaluated with linear probing by freezing ECG encoders and fine-tuning a linear layer on 1%, 10%, and 100% of labeled data from the six datasets. All Downstream tasks are evaluated using the average AUC across datasets, following the data splits from (Liu et al., 2024) (see Appendix A.6). Detailed setups are provided in Appendix A.7.
4.2 Evaluation with Mainstream eSSLs
We evaluate SuPreME against mainstream eSSL frameworks across 127 classes in six downstream ECG datasets, conducting linear probing with eSSL ECG encoders across varying data proportions to facilitate performance comparison. Table 1 demonstrates AUC results of SuPreME and eSSLs under different evaluation approaches.
@@ 4.2 Evaluation with Mainstream eSSLs
Table 1: Overall AUC performance of SuPreME and eSSLs, with ’Z’ for zero-shot and ’L’ for linear probing. Best results are bolded and second best gray-flagged.
Our results demonstrate that SuPreME achieves superior performance compared to traditional eSSL frameworks. With an overall zero-shot AUC of 77.20%, SuPreME outperforms most non-multimodal eSSL models, which require linear probing even with 1% (best: 59.46%) or 10% (best: 72.04%) of labeled data, showcasing its strong generalization capabilities and efficient utilization of pre-trained knowledge. Even without the CFN module, SuPreME remains highly competitive (in Appendix A.8) using only the pre-trained ECG encoder for linear probing. It consistently outperforms non-multimodal eSSL models across 1%, 10%, and 100% labeled data, and and achieves comparable performance to multimodal contrastive learning frameworks like MERL, despite not explicitly leveraging contrastive objectives. Notably, SuPreME achieves this with just 16 training epochs, compared to MERL’s 50 epochs.
Figure 8: Comparison of SuPreME (zero-shot learning) and eSSLs (linear probing with 1% data on the left and 10% data on the right) across downstream datasets.
Figure 8 presents framework performance across individual datasets. SuPreME’s advantage on the PTB-XL-Superclass dataset is minimal, likely due to the dataset’s simplicity, as it includes only 5 broad cardiac condition labels (e.g., NORM, STTC, MI), making it difficult to differentiate model performance. Additionally, all frameworks perform poorly on the PTB-XL-Form dataset, which focuses on 19 ECG waveform types that do not directly correspond to cardiac conditions, leading to ambiguous associations and reduced performance for all models.
To investigate SuPreME’s sensitivity to pre-training data, we evaluate its zero-shot performance across varying data proportions (in Figure 9). SuPreME’s performance improves steadily with more pre-training data, maintaining a significant edge over eSSL frameworks. Notably, with just 20% of pre-train data, SuPreME rivals or surpasses many eSSLs using 10% labeled data for linear probing. It also achieves comparable performance to multimodal eSSLs with only 60% of pre-train data, showcasing its efficiency and robust generalization, even in low-resource settings.
4.3 Evaluation of SuPreME Architecture
Beyond comparisons with eSSLs, we directly assess SuPreME’s zero-shot classification performance by varying its core modules, including the ECG backbone and CFN. Table 2 reports the results for SuPreME and its variants, where the ViT + CFN architecture achieves the highest average AUC of 77.20%, with particularly strong performance on PTB-XL-Rhythm and CSN.
@@ 4.3 Evaluation of SuPreME Architecture
Table 2: Zero-shot classification AUC performance of SuPreME and its variants on six downstream ECG datasets, with best results bolded.
Under the linear classification setup, ResNet outperforms ViT across most datasets, demonstrating its effectiveness in extracting essential features without contextual mechanisms. However, the CFN significantly improves performance, with SuPreME (ViT + CFN) achieving a notable boost in AUC, particularly on PTB-XL-Superclass (78.20%), PTB-XL-Rhythm (86.79%), and CSN (80.17%). Figure 10 highlights the strength of ViT’s attention mechanisms combined with CFN, which excels at capturing complex temporal and spatial dependencies in ECG signals.
By employing prompts to generate meaningful class queries, the CFN enables a more flexible and adaptable approach to classification, allowing flexible adaptation to diverse downstream class structures. Unlike linear classification, SuPreME dynamically aligns pre-trained knowledge with new cardiac conditions, eliminating the need for explicit class mapping between pre-training and downstream datasets and improving generalization.
@@
We further analyze performance on specific cardiac conditions in PTB-XL-Subclass (others in Appendix A.9), shown in Figure 11. SuPreME consistently achieves high AUC scores, particularly for critical conditions like LAFB/LPFB, CRBBB, CLBBB, IRBBB, and RVH (AUC > 90). In contrast, other variants show lower performance, especially for conditions requiring nuanced spatial and temporal patterns. ResNet + Linear performs competitively on simpler cases but struggles with complex conditions. ResNet + CFN exhibits significant drops, particularly for IMI, AVB, RAO/RAE, ISCA, and IRBBB, highlighting its limitations in effectively capturing intricate dependencies.
@@ 4.4 Ablation Analysis
Clinical Entity Mapping. To assess the impact of mapping labels to a standardized medical database, we compare SuPreME trained on the original 341 raw labels vs. our 295 deduplicated set (Table 3(a)). Notably, deduplication boosts the zero-shot AUC from 65.94% to 77.20%, which could be attributed to the removal of noisy or synonymous labels, thus clarifying the model’s representation of distinct cardiac conditions.
(a) Entity deduplication.
Deduplication Zero-shot AUC
Not Deduplicated 65.94
Deduplicated (Ours) 77.20
(b) ECG encoder backbones.
Backbone Zero-shot AUC
ResNet 64.96
ViT (Ours) 77.20
Table 3: Effects of pre-train dataset entity deduplication and different ECG encoder backbones.
ECG Encoder Backbone. To assess the ECG backbone impact, we replace the ViT-tiny with ResNet18 and evaluate zero-shot performance. ResNet18 yields a lower AUC of 64.96% (drops by 12.24%, Table 3(b)), indicating that ResNet struggles to capture longer-range ECG dependencies compared to ViT’s self-attention mechanism.
Clinical Text Encoder. For three different medical text encoders - BioClinicalBERT (Alsentzer et al., 2019), PubMedBERT (Gu et al., 2021), and MedCPT - MedCPT achieves the highest AUC, surpassing the others by 6.47% and 19.75% (Table 4(a)). This appears to stem from MedCPT’s contrastive objectives, which more effectively capture fine-grained sample similarities and differences than traditional masked language modeling.
(a) Language model encoders.
Language Model Zero-shot AUC
BioClinicalBERT 57.45
PubMedBERT 70.73
MedCPT (Ours) 77.20
(b) Cardiac Fusion Network.
Module Zero-shot AUC
w/o CFN (Linear) 71.23
CFN (Ours) 77.20
Table 4: Effects of different language models and Cardiac Fusion Network (CFN).
Cardiac Fusion Network Module. We compare CFN-based fusion with a simple linear projection (Table 4(b)). CFN lifts the zero-shot AUC from 71.23% to 77.20%, highlighting the benefits of cross-attention in capturing multimodal synergies between ECG signals and text queries.
Customized Cardiac Prompts. We evaluate three prompt strategies - GPT-4o Generated, CKEPE (detailed), and CKEPE (simplified). As shown in Table 5(a), the simplified CKEPE achieves the highest AUC, delivering at least a 12.5% improvement. This suggests that concise, clinically focused prompts help reduce noise and enhance alignment.
(a) Zero-shot query prompts.
Prompt Strategy Zero-shot AUC
GPT-4o Generated 61.30
CKEPE Detailed 64.70
CKEPE Simplified (Ours) 77.20
(b) Pre-train dropout ratios.
Dropout Ratio Zero-shot AUC
0.05 76.65
0.10 (Ours) 77.20
0.15 75.97
0.20 76.72
Table 5: Effects of different zero-shot query prompts and pre-train dropout ratios.
Dropout Ratio. During pre-training, we compare dropout rates of {0.05, 0.10, 0.15, 0.20}. In Table 5(b), a rate of 0.10 yields the best AUC (77.20%), likely striking a balance between regularization and signal retention.

### v2

@@ 3.3 Zero-shot Prompted Classification
To enable zero-shot classification on unseen cardiac conditions without fine-tuning, we construct concise, clinically meaningful prompts (e.g., left bundle branch block for LBBB) derived from SCP-ECG codes in each downstream dataset. These prompts form a dataset-specific query list aligned with the pre-training query space and serve as inputs to the textual modality of SuPreME.
We follow a simplified version of Clinical Knowledge-Enhanced Prompt Engineering (CKEPE) (Liu et al., 2024), where SCP-ECG codes are translated into discriminative phrases validated by UMLS and SNOMED CT. Unlike full CKEPE pipelines that retrieve verbose descriptions (e.g., a condition characterized by prolonged QRS complex... for LBBB), our approach promotes clarity and cross-modal fusion by avoiding redundant or overly detailed textual artifacts. These prompts are used exclusively during inference and remain fixed for all ECG samples within a dataset.
During zero-shot evaluation, ECG signals and textual prompts are encoded via the pre-trained encoders into $\mathbf{F}^{\text{eval}}_{\text{ECG}}$ and $\mathbf{F}^{\text{eval}}_{\text{Query}}$, then passed into the Cardiac Fusion Network (CFN). The CFN performs cross-modal attention to align features and outputs one logit per query-ECG pair. The final prediction scores are computed as:
$\mathbf{Pred}=\sigma(\text{CFN}(\mathbf{F}^{\text{eval}}_{\text{ECG}},\mathbf{F}^{\text{eval}}_{\text{Query}}))\in\mathbb{R}^{B\times M^{\prime}}$ (5)
This setup decouples the prediction space from any fixed label vocabulary, allowing the model to generalize to arbitrary diagnostic queries. The query list can vary across downstream datasets, and the classifier is query-agnostic, meaning no structural change is required when adapting to new tasks. Evaluation is conducted without fine-tuning using AUROC (AUC) per class and mean AUC across all prompts. Details about simplified-CKEPE, evaluation metrics are in Appendix A.5.
@@
Table 1: Performance of SuPreME and eSSLs, with ’Z’ for zero-shot and ’L’ for linear probing. Best results are bolded and second best gray-flagged.
@@
Table 2: Performance of SuPreME and its variants on downstream datasets. Best results are bolded.
Figure 4: Comparison of SuPreME (zero-shot) and eSSLs (linear probing with 1% data on the left and 10% data on the right) across downstream datasets.
From Scratch
ECG Only
Multimodal Learning
MERL 78.64 83.90 85.27 61.41 77.55 82.98 56.32 69.11 77.66 52.16 78.07 81.83 69.25 82.82 89.44 63.66 78.67 84.87
4 Experiments
4.1 Configuration and Settings
Clinical Entity Extraction. Following Section 3.1, we extract and normalize clinical entities from MIMIC-IV-ECG using Llama3.1-70B-Instruct [^2] Used offline for one-time inference only; not required during deployment. A large LLM ensures high-quality labels. (Appendix A.10.1) with structured prompts to ensure high-quality annotations. Entities are deduplicated via MedCPT embeddings (cosine similarity $>0.8$) and mapped to UMLS/SNOMED CT (average cosine similarity $>0.75$) [^3] Verified by cardiologists with 10+ years of experience; Appendix A.10.3.. Experiments are run on 8 NVIDIA A100-SMX4-80GB GPUs using vLLM Kwon et al. (2023). Statistics of extracted MIMIC-IV-ECG entities are in Appendix A.6.1.
Supervised ECG Pre-training. We use a 1D ViT-tiny encoder (patch size = 125, i.e., 0.25s) and a frozen MedCPT text encoder. Training employs AdamW (LR=$1\times 10^{-3}$, weight decay=$1\times 10^{-8}$) with cosine annealing ($T_{0}$=5000, $T_{\text{mult}}$=1, min LR=$1\times 10^{-8}$), for up to 50 epochs with early stopping (patience=10, best AUC at 16). Batch size is set to 256 on 4 NVIDIA A100-PCIE-40GB GPUs [^4] Compact pre-trained checkpoint (ViT-tiny + frozen MedCPT) runs on single GPU with $\geq$24GB memory, making it deployable in clinical or low-resource settings (Appendix A.10.2)..
Downstream Classification Task. SuPreME is evaluated on six unseen datasets (e.g., PTB-XL, CPSC-2018, and Chapman-Shaoxing-Ningbo) using dataset-specific prompts (Section 3.3). Ablation studies assess the impact of different ECG/text encoders and the CFN module as well as other key procedures. Mainstream eSSLs are benchmarked with linear probing by freezing ECG encoders and fine-tuning a linear layer on 1%, 10%, and 100% of labeled data from the six datasets. All tasks are evaluated by average AUC across classes and datasets, following the data splits in Appendix A.7.1. Hyperparameters are provided in Appendix A.7.2 and overlap analysis in Appendix A.6.2.
4.2 Evaluation with Mainstream eSSLs
We evaluate SuPreME against mainstream eSSL frameworks across 106 classes in six downstream ECG datasets, conducting linear probing with eSSL ECG encoders across varying data proportions to facilitate performance comparison. Table 1 demonstrates AUC results of SuPreME and eSSLs under different evaluation approaches.
Our results demonstrate that SuPreME achieves superior performance compared to traditional eSSL frameworks. With an overall zero-shot AUC of 77.20% (Details in Appendix A.8), SuPreME outperforms all non-multimodal eSSLs, which require linear probing even with 1% (best: 59.46%) or 10% (best: 72.04%) of labeled data, showcasing its strong generalization capabilities and efficient utilization of pre-trained knowledge. Even without the CFN module, SuPreME remains highly competitive (Table 3) using only the pre-trained ECG encoder for linear probing. Its overall performance consistently surpasses non-multimodal eSSL models across 1%, 10%, and 100% labeled data (also outperforms in 15/18 tasks with different datasets and labelled data portion), and achieves comparable performance to multimodal contrastive learning frameworks like MERL (ViT backbone with explicit contrastive objectives, more comparisons in Appendix A.10.4).
Since SuPreME’s ECG backbone is optimized jointly with the CFN rather than via an explicit contrastive loss, part of the diagnostic knowledge is encoded within cross-modal interactions. Linear probing on the ECG encoder alone thus cannot fully utilize the rich alignment captured during pre-training. This explains why the full SuPreME pipeline including query prompts and CFN achieves stronger zero-shot performance, even compared to linear probing with more labeled data. We highlight zero-shot performance as our main evaluation objective, as it aligns with real-world clinical settings where labeled ECG data is scarce and fine-tuning is often impractical.
Figure 4 presents framework performance across individual datasets. SuPreME’s advantage on the PTB-XL-Superclass dataset is minimal, likely due to the dataset’s simplicity, as it includes only 5 broad cardiac condition labels (e.g., NORM, STTC, MI), making it difficult to differentiate model performance. All frameworks perform poorly on the PTB-XL-Form dataset, which focuses on 19 ECG waveform types that do not directly correspond to cardiac conditions, leading to ambiguous associations and reduced performance for all models.
To investigate SuPreME’s sensitivity to pre-training scale, we evaluate its zero-shot performance under varying data proportions (Figure 5). SuPreME consistently improves with more data and maintains a clear advantage over non-multimodal eSSLs. Remarkably, with only 20% of pre-training data, SuPreME outperforms all non-multimodal eSSLs using 1% or 10% labeled data for linear probing, and matches the zero-shot performance of the multimodal baseline MERL trained on 100% of the pre-training data. Moreover, SuPreME’s zero-shot performance with just 20% of data also exceeds MERL’s linear probing result with 1% labels, highlighting its superior generalization and efficiency under limited supervision. Notably, SuPreME achieves these results with significantly fewer computational resources and shorter training time [^5] SuPreME: 4$\times$×A100-40GB GPUs for 16 epochs ($\sim$90 minutes); MERL: 8$\times$×A100-40GB GPUs for 50 epochs ($\geq$1 day). (Appendix A.10.2).
@@ 4.2 Evaluation with Mainstream eSSLs
LLM Size Zero-shot AUC
Llama3.1-8B-Instruct 72.89 $\pm$ 0.49
Llama3.1-70B-Instruct (Ours) 77.20 $\pm$ 0.21
Table 4: LLM for entity extraction.
Deduplication Zero-shot AUC
Not Deduplicated 65.94 $\pm$ 0.49
Deduplicated (Ours) 77.20 $\pm$ 0.21
Table 5: Entity deduplication.
Backbone Zero-shot AUC
ResNet 64.96 $\pm$ 0.20
ViT (Ours) 77.20 $\pm$ 0.21
Table 6: ECG backbone
encoders.
Language Model Zero-shot AUC
BioClinicalBERT 62.95 $\pm$ 0.53
PubMedBERT 62.51 $\pm$ 2.21
MedCPT (Ours) 77.20 $\pm$ 0.21
Table 7: Language model
encoders.
Module Zero-shot AUC
w/o CFN (Linear) 72.70 $\pm$ 0.42
CFN (Ours) 77.20 $\pm$ 0.21
Table 8: Cardiac Fusion
Network (CFN).
Prompt Strategy Zero-shot AUC
GPT-4o Generated 60.83 $\pm$ 0.26
CKEPE Detailed 69.16 $\pm$ 1.94
CKEPE Simplified (Ours) 77.20 $\pm$ 0.21
Table 9: Zero-shot cardiac query prompts.
Dropout Ratio Zero-shot AUC
0.05 75.98 $\pm$ 0.56
0.10 (Ours) 77.20 $\pm$ 0.21
0.15 75.63 $\pm$ 0.63
Table 10: Pre-training
dropout ratios.
4.3 Evaluation of SuPreME Architecture
Beyond comparisons with eSSLs, we assess the contribution of core components in SuPreME by varying its core modules, including the ECG backbone (ResNet vs. ViT) and (Linear vs. CFN) shown in Table 2 and Figure 6. Overall, SuPreME (ViT + CFN) achieves the highest average AUC of 77.20%, with strong results on PTB-XL-Rhythm (86.79%) and CSN (80.17%), demonstrating the effectiveness of cross-modal fusion for temporally and spatially complex signals.
Under linear classification, ResNet outperforms ViT across most datasets, reflecting its inductive bias toward local feature extraction. However, once CFN is introduced, ViT significantly benefits from its attention mechanisms and structured prompts, outperforming all other variants. This suggests that ViT’s global receptive field aligns well with the query-driven fusion in CFN, while ResNet’s local filters are less suited for attending over sparse textual queries.
Notably, the performance of ResNet + CFN is lower than ResNet + Linear across several datasets. We attribute this to a mismatch between ResNet’s hierarchical, spatially localized features and CFN’s attention-based fusion, which benefits more from globally contextualized inputs like ViT. CFN is designed to interpret semantically aligned queries over long-range dependencies, an area where ResNet lacks representational flexibility. This highlights the importance of matching the backbone’s encoding characteristics with the fusion strategy. Details are in Appendix A.10.5.
Figure 7 further analyzes SuPreME’s performance on individual cardiac conditions in PTB-XL-Subclass (Complete results in Appendix A.9). SuPreME consistently achieves high AUCs (many $>90$), especially for nuanced conditions like LAFB/LPFB, CRBBB, CLBBB, and RVH, where both query semantics and signal patterns must be integrated. In contrast, ResNet + CFN underperforms in complex arrhythmias (e.g., ISCA, IRBBB), reinforcing our insight that strong multimodal fusion requires compatible encoders.
Unlike linear classifiers with fixed output dimensions, CFN enables flexible prompt-driven classification, aligning query semantics with signal patterns in a shared latent space as specified in Section 3.2, improving generalization to novel conditions without fine-tuning and demonstrates its strength especially when paired with ViT.
@@ 4.4 Ablation Analysis
Entity Extraction Model. Using Llama3.1-70B-Instruct for NER improves zero-shot AUC by 4.31% over the 8B variant (Table 4), reflecting better supervision quality. This step is offline and serves to build a high-quality labeled dataset, not for deployment.
Clinical Entity Mapping. Mapping to a standardized 295-term vocabulary improves zero-shot AUC from 65.94% to 77.20% (Table 5), likely by removing noise and resolving label redundancy to better represent distinct cardiac conditions.
ECG Encoder Backbone. Replacing ViT-tiny with ResNet18 drops zero-shot AUC by 12.24% (Table 6), suggesting ResNet is less effective at modeling long-range ECG dependencies than ViT (Appendix A.10.5).
Clinical Text Encoder. Among BioClinicalBERT, PubMedBERT, and MedCPT, the latter achieves the highest AUC, outperforming the others by over 14.25% (Table 7), likely due to its contrastive training objective, which better captures fine-grained clinical distinctions.
Cardiac Fusion Network Module. We compare CFN-based fusion with a simple linear projection (Table 8). CFN lifts the zero-shot AUC from 72.70% to 77.20%, highlighting the benefits of cross-attention in capturing multimodal synergies between ECG signals and text queries.
Customized Cardiac Prompts. Among three strategies (GPT-4o, detailed CKEPE, and simplified CKEPE), the simplified CKEPE achieves the best AUC (Table 9), with $\geq$ 8.04% improvement, suggesting that concise, clinically focused prompts enhance alignment and reduce noise.
Dropout Ratio. In pre-training, we compare dropout rates of {0.05, 0.10, 0.15}. A rate of 0.10 yields the best AUC, striking a balance between regularization and signal retention (Table 10).